OpenStudy (anonymous):

Can you help me integrate this function: ln(e*x) * 1/x. between e and 1.

hartnn (hartnn):

use the log property first \(\ln AB = \ln A + \ln B\)

hartnn (hartnn):

ln e =1

hartnn (hartnn):

then you need to do a u substitution

OpenStudy (anonymous):

1/x + 1/x*ln(x) is that what it is so far?

hartnn (hartnn):

yeah, but keep it as \((1+\ln x)/x\) now do a substitution u = 1+ln x du = ... ?

OpenStudy (anonymous):

du/dx = 1/x so integral .... du = integral ... 1/x dx

hartnn (hartnn):

your new integral in terms of "u" will be ??

OpenStudy (anonymous):

between 2 and 1?

hartnn (hartnn):

yes so we have  got this ?

OpenStudy (anonymous):

Yup, thanks, I pretty much understand. So whenever I do a substitution integration, all I end up integrating is u?

hartnn (hartnn):

not everytime! in this case 1+ log x was plugged in as 'u' thats why numerator has u

OpenStudy (anonymous):

Ok, and how do I integrate 1+lnx ?

hartnn (hartnn):

what ? we did change that to "u" , isn't it ?

hartnn (hartnn):